tangents and normals
find the equation of the tangent line to the curve y=x^2 – 2x + 7 which is parallel to the line 2x – y + 9 = 0

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

y = x^2 - 2x + 9
Slope of tangent: y’ = 2
y' = 2x - 2
2x - 2 = 2
x = 2
y = 9
Equation:
(y - 9) = 2(x-2)